### v1

# Probing Non-Thermal Processes in Stellar Flares on AU Mic

Abstract
The role of non-thermal proton energy transportation during solar and stellar flares is largely unknown; a better understanding of this physical process will allow us to rectify longstanding deficiencies in flare models. One way to detect the presence of non-thermal protons during flares is through the Orrall-Zirker (OZ) effect whereby an enhanced red wing appears in hydrogen emission lines (e.g., Lyman- at 1215.67 Å). We analyze archival Hubble Space Telescope/Cosmic Origins Spectrograph G130M (1060 - 1360 Å) observations of the young M dwarf, AU Mic to search for evidence of OZ effect during the impulsive phase of six stellar flares with erg. While we found non-detections of the OZ effect, we note there is a pronounced blue enhancement in several C II and C III emission lines during one of the high-energy flares. We propose that either filament eruptions or chromospheric evaporation could be the mechanism driving this observed blue enhancement. We compare the far-ultraviolet (FUV) spectra to 1D radiative-hydrodynamic stellar flare models, which are unable to reproduce the blue enhancement and broadening in these cool flare lines. By completing a line-by-line analysis of the FUV spectrum of AU Mic, we provide further constraints on the physical mechanisms producing stellar flares on M dwarfs.
1 Introduction
Our knowledge of flare-accelerated particles during magnetic reconnection events on other stars has significant gaps. The mechanism behind particle acceleration in flares, the resulting particle energy distributions, and in particular the role of non-thermal protons are poorly understood and are lacking in current flare models (Kowalski, 2024). Proton beams may constitute up to or exceeding half of the energy during particle acceleration in solar flares (Emslie et al., 2012). If this is the case, then our current flare models are missing an important constituent of the energy transport through the solar/stellar atmosphere (Kerr et al., 2023). Non-thermal protons are thought to deposit their energy deeper in the atmosphere than non-thermal electrons, thus possibly providing an explanation for the enhancements in the optical continuum, commonly referred to as white-light, that have a photospheric origin (Sadykov et al., 2024; Kerr & Fletcher, 2014; Watanabe et al., 2013; Neidig, 1989; Procházka et al., 2018).
Non-thermal electrons cannot account for a photospheric origin of white-light flares because they cannot reach deep enough into the atmosphere to cause this heating (Watanabe et al., 2013). To determine the cause of white-light flares and other responses to deep energy deposition, we seek an accurate picture of the non-thermal proton acceleration process. Constraining the presence of low-energy protons ( MeV) among the accelerated particle population in turn constrains the total energies available to explain the white-light radiation and other flare emission. An empirical re-examination of the evidence for these low-energy protons is timely because there has been significant progress in producing power-law spectra of protons extending down to low energies during reconnection (Yin et al., 2024) and in modeling the warm-target transport of low-energy protons from the corona to the chromosphere (Allred et al., 2020). Furthermore, re-examination of gamma-ray images of large solar flares now suggests the sources of non-thermal electron and proton impact with the chromosphere to be cospatial (Battaglia & Krucker, 2025), thus transforming decades of understanding (Hudson, 2016).
One possible observational signature of non-thermal protons accelerated during stellar flare events is the Orrall-Zirker (OZ) effect (Orrall & Zirker, 1976). The OZ effect occurs when non-thermal protons are accelerated downwards in the corona to the mid-upper chromosphere. In the chromosphere, the non-thermal protons charge exchange with ambient hydrogen atoms, thus producing a population of non-thermal energetic neutral atoms (ENAs). These ENAs subsequently go through spontaneous radiative de-excitation and emit Doppler redshifted photons in the red wings of hydrogen spectral lines (Figure 4 in Kerr et al., 2023). Tentative evidence of the OZ effect has been seen in a single star — AU Microscopii (AU Mic; Woodgate et al., 1992). AU Mic is a Myr (Bell et al., 2015) M-dwarf (M1Ve; ; Keenan & McNeil, 1989; Plavchan et al., 2020; Gallenne et al., 2022). AU Mic is known for its enhanced stellar activity due to its young age, including a high flaring rate (Feinstein et al., 2022; Gilbert et al., 2022; Tristan et al., 2023), short rotation period ( days; Plavchan et al., 2020), and strong magnetic field (Donati et al., 2023). AU Mic has been the target for several flare-driven observational campaigns (Tristan et al., 2023; Paudel et al., 2024; Tristan et al., 2025) and is a standard for modeling M-dwarf flares due to its close proximity ( pc; Gaia Collaboration et al., 2021) and high signal-to-noise observations.
Orrall & Zirker (1976) predicted a signature of non-thermal proton beams accelerated during flares could be observed as an increase in the red wing of Lyman- (Ly; 1215.67 Å). Kerr et al. (2023) defined the expected wavelength range to detect the OZ effect as 5.6 - 17.7 Å redward (1380-4365 km s-1) of Ly (see also Canfield & Chang, 1985). In 1992, Woodgate et al. (1992) presented a tentative detection of the OZ effect on AU Mic through an observed red enhancement in Ly using the Goddard High Resolution Spectrograph on the Hubble Space Telescope (HST). The detection lasted for 3 s at the flare start, which is in agreement with the requirement for a relatively large area of undisturbed atmosphere. To date, this is still the only tentative detection of the OZ effect on another star (Kerr et al., 2023; Robinson et al., 2001). This tentative detection highlights that HST could detect the OZ effect here if a sufficient amount of non-thermal protons are accelerated during a flare. Feinstein et al. (2022) performed a preliminary search in recent HST/Cosmic Origins Spectrograph (COS) observations and did not detect the presence of the OZ effect.
In addition to its activity, AU Mic has been a target of observational campaigns due to the presence of two transiting exoplanets (Plavchan et al., 2020; Martioli et al., 2021) and its debris disk (Liu, 2004; Lawson et al., 2023). AU Mic b is a Neptune-like exoplanet () on a close orbit, with a semi-major axis au. AU Mic c is slightly smaller ( ) and farther than AU Mic b, with a semi-major axis of au (Martioli et al., 2021). Due to the close proximity of both planets, their evolution has the potential to be shaped by AU Mic’s frequent stellar flares (e.g., Feinstein et al., 2022; do Amaral et al., 2025). For example, M-dwarf stars’ frequent and intense flaring may be a source of atmospheric stripping of exoplanets Lammer et al. (2007), with exoplanet atmospheric loss being predicted to be partially driven by M-dwarf flaring activity (Neves-Ribeiro et al., 2022). The potential of later atmospheric recovery is debated (Tilley et al., 2019). On the contrary, frequent flaring may jump-start the formation of organic molecules and prebiotic chemistry necessary for the evolution of life (Dransfield et al., 2024); simulations of superflares on the young Sun suggest they may have been instrumental in forming the molecules essential for life (Airapetian et al., 2016).
Beyond the AU Mic system, a complete understanding of the underlying physics driving magnetic reconnection events and the resulting observable stellar flares will provide a comprehensive understanding of the environments in which most planets evolve. M-dwarf stars such as AU Mic are known for their frequent and intense flaring throughout and especially early in their lives (Feinstein et al., 2020b; Crowley et al., 2024). Stellar CMEs, often accompanied by flares, can potentially be associated with spectral line shifts during a stellar flare, as shown in Ikuta & Shibata (2024). Due to the close-in habitable zones around M-dwarfs, atmospheric erosion of potentially habitable planets can be brought about by stellar CMEs from M-dwarfs more easily than CMEs from other spectral types of stars. It is crucial to understand the frequency and time evolution of stellar CMEs from M-dwarfs in order to determine their impact on atmospheric loss and evolution (Khodachenko et al., 2007; Lammer et al., 2007). The current models of stellar flares are not conclusive or complete enough to give accurate insight into the effects of stellar flares on exoplanet atmospheres. Bridging the current gaps in how flares operate will allow an understanding of exoplanet atmospheric responses.
In this work, we present a reanalysis of three HST/COS visits of AU Mic (Feinstein et al., 2022, 2024) to search for evidence of non-thermal proton beams. We present newly identified flares in the third visit. This work is laid out as follows. We describe the HST observations, data reduction, and emission line identification in Section 2. We identify and characterize new stellar flares in Section 3 and present the first far-ultraviolet (FUV) flare-frequency distribution (FFD) for AU Mic. In Section 4, we present the results of our search for evidence of the OZ effect and quantify an identified blue enhancement in several C II and C III emission lines. In Section 5 we speculate further on the cause of the identified carbon enhancement and discuss the implications of this work on stellar flare models. We conclude in Section 6.
2 Observations
We observed three visits of AU Mic with HST/COS G130M (Å; Å) under GO 16164 (PI Cauley). Visit 1 was executed on 2021-05-28 from 03:33:31 — 15:50:43, Visit 2 was executed on 2021-09-23 from 15:06:14 — 22:02:08, and Visit 3 was executed on 2022-10-09 from 12:38:55 — 19:33:40. The instrument configuration for all three visits is identical. All visits have a total exposure duration of 2085 seconds for the first orbit and an exposure duration of 2471 seconds for the subsequent orbits.
COS has a unique time-tag mode, which tracks the time at which each photon hits the detector. Through this mode, we are able to generate time-tagged spectra and light curves. We used the costools.splittag111https://github.com/spacetelescope/costools/tree/master/costools and CalCOS222https://github.com/spacetelescope/calcos/tree/master/calcos pipelines to process the data into wavelength-calibrated spectra and analyzed the spectrum from each corrtag file, using these spectra to create spectroscopic light curves. We generated light curves at a 30-second cadence, which is sufficient for studying the morphology of flares. While we explored light curves at a higher cadence, the signal-to-noise (SNR) of individual emission lines became too low. Thus, the 30-second light curves provided a balance between a high time resolution and sufficient SNR to analyze specific emission lines. We note that the wavelength calibration region for the fourth orbit for Visit 3 was affected by a hot spot, a transient region of high detector background. The standard costools and CalCOS pipelines were unable to fully reduce this visit while the hot spot was present. As such, in order to uniformly reduce this orbit with the others during Visit 3, we masked the hot spot in the rawtag file. This hot spot occupies a region of the detector between 6850 RAWX 6950 and 570 RAWY 610.
AU Mic has a very high SNR FUV spectrum, allowing for the identification of over 100 individual emission lines (Feinstein et al., 2022). For the purposes of our investigation, we used several emission lines: C III six-tuplet ( Å), Ly , and C II doublet ( Å).
3 The FUV Stellar Flares of AU Mic
In this work, we present new stellar flares which occurred during Visit 3. We followed the identification technique described in Feinstein et al. (2022). Feinstein et al. (2022) identified flares in Visits 1 and 2 by identifying outliers in the Si III ( = 1294.55 Å) and C III ( = 1175.95 Å) emission line light curves, presented in Figure 1. Flares were identified by their well-known time-dependent behavior of a sharp initial increase followed by an exponential decay phase. We identify an additional seven flares in Visit 3; we follow the same naming convention and introduce flares N - T.
We uniformly calculated the energies and equivalent durations for flares A - T, where flares A - M occurred in visits 1 and 2. We derive the flare energy using the following prescription:
where is the distance to the star, is the start time of the flare defined as the time point the flux increases by from the quiescent flux, is the stop time of the flare defined as the time point the flux gets within of the average quiescent flux, , and is the total flux during the defined start and stop times. We adopted (Gaia Collaboration et al., 2021). To calculate the equivalent duration (ED), first described by Gershberg (1972), we used the following prescription:
We calculated the energies and durations of the new flares in the “white light” (broadband) wavelength-integrated flux ( Å), Si III wavelength-integrated line flux, and C III wavelength-integrated line flux. For the white light flux, we masked the Å due to the lower signal-to-noise at these wavelengths. We repeated this process for the entire sample of flares. We note that Flare R (Figure 1) has an extended decay, which could be due to the presence of lower energy flares we are unable to resolve. This event could be influencing our ability to accurately calculate the true and for Flare S as the flux does not return to the pre-flare baseline before Flare S begins; this can be seen in both the white light and C III light curves and flare properties. The measured properties of all flares are presented in Table 2.
3.1 Optical Flares on AU Mic
AU Mic was observed with the Transiting Exoplanet Survey Satellite, (TESS; Ricker et al., 2015), in Sectors 1 (August 2018) and 27 (July 2020). Gilbert et al. (2022) and Tristan et al. (2025) have both presented independent analyses of the TESS flares on AU Mic. We present a reanalysis using a third independent technique.
We downloaded the TESS 2-minute cadence observations through lightkurve (Lightkurve Collaboration et al., 2018). We use the PDC_SAP flux from the SPOC-processed light curves (Jenkins et al., 2010). We identify flares using the convolutional neural network (CNN) model stella (Feinstein et al., 2020a, b). stella assigns each data point a probability that it is (1) or is not (0) part of a flare. We choose this method as the CNN model is able to identify flares without removing rotational photometric variability. We considered all flares with an assigned probability .
We calculated the flare energies and EDs using Equations 1 and 2. We identified 54 flares in Sector 1 and 55 flares in Sector 27, for a total of 109 flares in 48.7 days of observations. This results in an optical flare rate of 2.24 flares day-1. Our sample of TESS flares has erg. Our recovered number of flares is slightly higher than but still consistent with previous works (Gilbert et al., 2022; Tristan et al., 2025). We note that additional TESS observations of AU Mic are expected in Sector 95 (starting August 2025).
3.2 FUV and Optical Flare Frequency Distributions
We compare the FUV and optical flare frequency distributions (FFD) of AU Mic. We derive our FUV FFD using the white light HST, Si III, and C III flare energies derived in this work. We look at the FFD of multiple spectral lines as this may provide insights into how the emission processes may change over the range of observed flares. We derive our optical FFD using observations from the TESS observations. All resulting FFDs are presented in Figure 2. We fit two distributions to describe the FFD. First, we fit a truncated power law of the form
where is the energy, is a normalization factor, is the slope, is the energy at which the function turns over, is an offset factor, and is the flare rate. We fit this form to characterize . We use the non-linear least squares model optimization implemented with scipy.optimize.curve_fit to fit for each variable. We repeat this process for all FFDs. We fit all available energy bins, with the exception of the highest two energy bins in the HST WL and C III FFDs. We choose to ignore these bins as the high-energy regime of the FFD is sparsely sampled due to the limited baseline of the HST observations.
Additionally, we fit a piecewise function of two power-laws to measure the slope, , of the distribution. We choose to use this as the descriptor of the slope, rather than in the truncated power-law, as measuring is a more common practice when describing FFDs (e.g., Ilin et al., 2019, 2021). We defined our piecewise function as:
where is the slope and is the y-intercept. We present the best-fit values for , , and in Table 1. Due to the large uncertainties in for the HST WL and Si III FFDs, we fit these distributions as a single power-law.
FFD | [erg] | ||
TESS | |||
HST | — | ||
WL | — | ||
HST | — | ||
Si III | |||
HST | |||
C III |
We compare the measured TESS FFD to the average FFD of stars of similar ages and to AU Mic presented in Feinstein et al. (2024). We find that is consistent to within of the FFD slope for stars with K and ages Myr (Feinstein et al., 2024). We find that is inconsistent with the measured FFD slope of stars of similar temperatures and age to AU Mic by . This discrepancy could be due to the high SNR observations of AU Mic, which result in a more complete FFD, or could be an indication that AU Mic is particularly active for its age. A third alternative scenario is that the flares of AU Mic are driven by star-planet interactions with its two short-period gas-giant planets (Ilin & Poppenhaeger, 2022; Feinstein et al., 2024).
Loyd et al. (2018) presented an analysis of inactive and active M dwarfs that were observed as part of the Measurements of the Ultraviolet Spectral Characteristics of Low-mass Exoplanetary Systems (MUSCLES) Treasury survey (France et al., 2016). This study included six “inactive” M dwarfs observed with COS/G130M and four “active” M dwarfs STIS/E140M, including AU Mic. Loyd et al. (2018) measured the FUV flux within specific integrated regions and fit a single power-law with a slope of for active M-dwarfs. We note that the inactive M dwarfs had a flare sample size similar to that presented here. The shallower FFD slope measured here could indicate that AU Mic is more active than the cumulative sample of M-dwarfs. More observations of AU Mic in the FUV are required to properly sample the high-energy regime of the FFD.
4 Searching Flare-Affiliated for Non-Thermal Processes
A detailed analysis of the behavior of individual emission lines during flare events can yield greater insight into the overall physical mechanisms driving stellar flares. It is only with high signal-to-noise observations of M dwarfs during large flares, such as those presented here, can we probe these effects.
4.1 The Orrall-Zirker Effect
The Ly emission line was masked in the observations to avoid detector saturation from 1206 - 1220 Å. Even with this mask, the spectral range is sufficient to search for evidence of this effect in the HST/COS data set. Should a non-thermal proton beam be present during these flares, we would expect to see the red wing emission enhancement during the early impulsive phase of a flare (Woodgate et al., 1992), when much of the atmosphere among the flaring areas is neutral. Thus, we calculated spectroscopic light curves of the red wing (1220 - 1233.37 Å) and blue wing (1197.97 - 1206 Å) of the Ly emission line to detect the OZ effect. Within the blue wing, we masked S III ( = 1200.97, 1206.53, 1207.51 Å), as identified using the HST/STIS spectrum of AU Mic presented in Rockcliffe et al. (2023a). This left a sufficient baseline to determine if the blue wing responded asymmetrically to the red wing in the observed flares. This analysis was repeated for six of the highest-energy flares in our data set: Flares B, J, K, M, R, and T .
We present example time-dependent spectra of the Ly wings during Flare B in Figure 3. We show the spectra of AU Mic pre-flare (yellow), the flare impulsive phase (light green), the flare peak (green), and the secondary flare peak (blue). The maximum change in the blue wing flux is consistent with the maximum change in the red wing flux to within . Thus, we find there is no strong evidence of the OZ effect during this flare event. We repeated this analysis for Flares J, K, M, R, and T, which are all consistent with the findings for Flare B.
4.2 Observed Carbon Blue Enhancement
We describe a blue enhancement in Flare B in two emission lines: C III ( Å) and C II ( Å). The centroids of single Gaussian fits are considered as the overall blueshift for the lines. For Flare B, we calculated a maximum blueshift of km s-1 in C III and km s-1 in C II. Despite the lines being complicated mixtures of several transitions, a single Gaussian fit to each resolved component is sufficient to elucidate general velocities.
Flare B is a large triple-peaked flare event and the longest-lived event observed in the FUV. We present the blue wing, core, and red wing light curves of C II and C III in Figures 4 and 5, respectively. Both figures show that the blue wings of C II and C III appear to respond mainly to the first peak in Flare B, which is further evidence that the second peak in Flare B is a second flare due to its different spectral responses. In the second peak, the blue enhancement of the C III and C II spectral lines is not as pronounced as in the first peak. The line response in the second peak seems to be similar to other lines during the flare. In fact, we see in C II and C III that the second flare in the Flare B event has a stronger response in the red wing than the first flare. This highlights the unique emission line response for sequential flares. This response was analyzed in other C emission lines, but these emission lines lacked the signal-to-noise needed to make a clear observation of the spectral response of the blue and red wings throughout Flare B.
4.3 Modeling Chromospheric Evaporation
Chromospheric evaporation occurs when non-thermal electrons stream along magnetic field lines and impact the chromosphere. Once these non-thermal electrons arrive at the chromosphere, they deposit their energy into the plasma, heating the plasma and increasing pressure. The heated plasma is accelerated upwards, resulting in chromospheric evaporation Fisher et al. (1985). This phenomenon is observed in the Sun. Observational evidence of chromospheric evaporation includes blue-shifted spectral lines during the impulsive phase of the flare (Milligan & Dennis, 2009; Antonucci et al., 1984; Shibata et al., 1995). If the blue-enhancement seen in C II and C III is due to chromospheric evaporation, this would imply that the gas emitted at K is being accelerated upwards through the flare loop. The blueshift would be seen at the foot points of the soft X-ray (SXR) loop depending on its temperature. Here, we aim to determine if models of solar-like chromospheric evaporation can be used to explain the observed C II blue enhancement.
We used a 1D radiative-hydrodynamic M-dwarf stellar flare model from the grid described in Kowalski et al. (2024) to model the spectrum of C II during Flare B. We generated a flare atmosphere model with the following assumptions: (i) peak injected electron beam energy flux density of [erg cm-2 s-1], (ii) a low energy cutoff of 37 keV, and (iii) a power law index of = 2.5. We chose this model because it is able to reproduce M dwarf optical hydrogen lines well when combined with a model for the continuum (Kowalski, 2022; Kowalski et al., 2025). The chosen model also produces upflows of several hundred km s-1 that could inform our understanding of mass motions during the flare events. This model is a higher energy extension of the F11 solar flare model in Allred et al. (2005), which produces prominent blueshifts in cool temperature ( K) flare lines.
We modeled individual time steps between the impulsive and decay phases of Flare B using the RH radiative transfer code from Uitenbroek (2001). We evaluated the model for Flare B between seconds at s time-intervals. Note that the model timescale is different than the timescale of Flare B, which lasts seconds (2). The models in the grid represent a superposition of short bursts; nonetheless analysis of individual time steps is necessary to understand how the atmosphere responds in a spatially unresolved observation. To synthesize the C II emission line, we use the model atom from (Rathore et al., 2015). A handful of time steps required modification to the number of depth points () included in the model. Initially, the model had . We had to reduce to 180 in most cases, excluding heights in the upper corona, to achieve convergence. As the populations of C II are very small in these hot regions, truncating the atmosphere in this way does not affect our results. Additionally, time steps from s required . We concluded this would be an insufficient height representation for an accurate stellar flare model. Therefore, we do not include models at s for C II in our analysis. All models are calculated at 5 emergent rays ( 18.19∘, 38.74∘, 60.00∘, 76.70∘, and 87.13∘), and are summed using Gaussian quadrature to give an emergent radiative flux spectrum.
The model comparisons are shown in Figure 6. While every model has limitations, we can draw several conclusions concerning our observation of the response of the C II emission lines during Flare B from the Kowalski et al. (2024) model. First, we find that the indirect emerging ray ( = 87.13∘) contributes to line broadening, while the direct emerging ray ( = 18.19∘) contributes to the spectral line strength. However, none of the rays or the flux account for the discrepancy between the blue and red enhancements of the C II spectral line; there is no significant redshift or blueshift in the model for C II.
The source of symmetric broadening in non-hydrogenic lines in solar flares is yet not well understood (Kerr, 2023). We follow Zhu et al. (2019) and multiply the Stark-B333Sahal-Bréchot, S., Dimitrijević, M.S., Moreau N., 2025. STARK-B database, [online],444http://stark-b.obspm.fr [June 9, 2025]. Observatory of Paris, LERMA and Astronomical Observatory of Belgrade damping factor () for the C II lines by a factor of 30 (Mahmoudi et al., 2004, 2005; Larbi-Terzi et al., 2012). The model with 30 for C II (Figure 4 better accounts for the shape of the spectral line throughout the flare. However, the model still does not explain the discrepancy between the blue and red enhancements of the C II spectral line. Quantifying the badness of fit for the models, we find that the RH calculation 30 had a reduced , throughout all the times modeled in Figure 6, and we find that the RH calculation without the added factor of 30 had a reduced , . For the C II models without the added factor of 30, the general line broadening is not well accounted for, as is apparent in Figure 6. There is also a poor representation of the intensity of the two different lines in the C II doublet, with the bluer C II line ( = 1334.53 Å) being less intense than the redder C II line ( = 1335.66 Å) in the observations but less so in the model.
5 Discussion
Woodgate et al. (1992) remains the only tentative detection of the OZ effect to date. Non-thermal proton beams may still be present in flares on AU Mic; however, the new observations exclude copious fluxes of low-energy protons ( MeV; Kerr et al., 2023) that cause the OZ effect. Higher energy protons in solar/stellar flares could potentially be detected with -ray and X-ray observations (Kerr et al., 2020; Battaglia & Krucker, 2025).
We compare the observed C II blueshifts against a well-established chromospheric evaporation model from Kowalski et al. (2024); we found this model did not account for the observed asymmetries, which has led us to propose a second possible mechanism. Filament eruptions often accompany solar flares; the same could be true here. Filaments form in the lower atmosphere and contain cool, dense plasma suspended by magnetic fields. Destabilization and filament eruptions could occur as a result of the magnetic reconnection event and vice versa. The filament eruption happens in the upper threaded loop during this process (Moore et al., 2001). After the low-lying filament erupts, pulling the flare loop up through the atmosphere, the reconnection event happens underneath (Shibata et al., 1995).
In Ikuta & Shibata (2024), a one-dimensional free-fall model and a pseudo-2D MHD model were used to infer the direction of the ejection, changes in height, and the possible development of a magnetic loop during a flare on EK Draconis. A significant redshift of the H spectral line was predicted; this magnitude of redshift is not present in the C II or C III spectral lines during Flare B. We speculate that the blue enhancement observed here for C III and C II could be explained by emission coming from the expansion of the legs of the filament. This expansion would produce an initial blueshift, followed by a redshift as the filament expands. The phenomenon was reported for the H line during a superflare on EK Draconis, in Ikuta & Shibata (2024) the H line variance was modeled during a flare, and it was found that filament eruptions could explain the variance observed. The blue enhancement caused by the filament eruptions has a particular velocity profile starting at velocities of a few tens of km/s up to hundreds of km/s back to zero, and then there is a redshift later on; this process is illustrated in Figure 10 of Ikuta & Shibata (2024). In Ikuta & Shibata (2024), there was an initial acceleration of over 500 km/s for the H spectral line, decelerating to -200 km/s. During Flare B, there was an initial similar blue acceleration of C II and C III, in the range of a few hundred km/s.
Current abilities to model broadening of Mg II for solar flares fall drastically short (Zhu et al., 2019). Our research highlights that a similar problem potentially exists for C II. Zhu et al. (2019) introduced a Stark width factor of 30. We explore the role of this for C II, shown in Figure 6. By multiplying by a factor of 30 for the C II resonance lines, the model is better able to describe the broadening throughout the flare; this suggests that an unknown source of atomic line broadening may also explain the C II widths in stellar flares. The enhanced symmetric broadening extends far into the wings (Figure 6, s). Determining the physical origin of the missing amounts of the broadening would thus improve the interpretation of asymmetric wings due to mass motions. In future work, we plan to explore the role of this parameter in more detail while including the C III lines (which are not included in the Rathore et al. 2015 model atom) in the calculations.
5.1 FUV Flare Fitting
The flare energies and equivalent durations presented here were derived from the light curve. Other work have reported these values derived from a best-fit model. We attempted to model these FUV flares with the open-source code Llamaradas Estelares (Tovar Mendoza et al., 2022), an analytical flare model that was optimized for optical flares. We present an example fit to Flare R in Figure 7. The Llamaradas Estelares flare model was unable to capture the decay phases of several flares, including Flare R, accurately. In particular, the model was unable to account for the extended tail of Flare R, which could be due to subsequent sympathetic flares or persistent emission lines during the decay phase. We emphasize exercising caution in future work when using this model to calculate flare energy and duration, as it has not consistently provided an accurate fit in this set of flares. A more flexible model is needed for modeling FUV flares.
6 Conclusions and Future Work
In this paper, we present a reanalysis of Hubble Space Telescope/Cosmic Origins Spectrograph FUV observations of the 25 Myr planet-host AU Mic. Our analysis focused on searching for evidence of non-thermal processes across the 20 identified flares in these observations. We conclude the following.
We present the first comparison of the FUV and optical flare frequency distributions (FFDs) of AU Mic (Section 3.2). The differences in the FFDs suggest that the average energy budget of the flares dominates in the FUV over the optical, which is consistent with our understanding of flare temperatures.
We find no evidence of the OZ effect in the high-energy flares analyzed (Section 4); Woodgate et al. (1992) remains the only tentative detection of the OZ effect on another star to-date. Our non-detections provide insight into the frequency of proton beams in M-dwarf flares and can be used to enhance our models of such events, which currently lack energy transport from non-thermal protons.
We find evidence of blue wing enhancements in two spectral lines, C II ( Å) and C III ( Å), during Flare B (Section 4.2). We propose these enhancements may be caused by filament eruptions or chromospheric evaporation. However, we note that the observations may favor filament eruptions, as we tested the observed blueshifts against the well-established chromospheric evaporation model presented in Kowalski et al. (2024).
We modeled the line strength and shape of C II using the electron beam heating grid provided by Kowalski et al. (2024) and found that under the assumptions of the current grid, we are unable to accurately recreate the spectral line strength and broadening of C II (Section 4.3). We find this broadening is better explained by multiplying the model by the Stark-B damping factor, .
Our work highlights the lack of understanding of the response of specific emission lines during M dwarf flares. The results of this study will improve our modeling assumptions of M dwarf flares and may have implications for modeling exoplanet atmospheric responses to stellar flares.
6.1 Future Work
Additional archival and future observations of AU Mic are important for understanding the occurrence of the OZ effect (e.g., HST GO 17613). AU Mic was observed with HST/STIS observations in 2020-2021 (HST-GO-15836; PI: Newton). This instrument configuration is centered on Ly (Rockcliffe et al., 2023b). These observations could potentially show the presence of the OZ effect and should be analyzed in future work.
Here, we focused on using a well-established chromospheric evaporation model from Kowalski et al. (2024) to model the C II line. The work presented here could be extended to modeling C III. Additional insights from other carbon emission lines could help solidify whether the blue wing enhancements seen during flare events are driven by chromospheric evaporation or an alternative mechanism. Additional, simultaneous observations in the soft X-ray and other ultraviolet wavelengths could also be used to better understand this phenomenon.
In solar flares, we expect the transition from blueshifts to redshifts to occur at about 1 MK, according to Milligan & Dennis (2009). However, in these stellar flares, we observe that this does not hold. We are observing blueshifts in lines that are significantly cooler than 1 MK, C II (log K; Hawley et al., 2003) and C III (log K; Hawley et al., 2003) in Flare B. The chromospheric evaporation scenario that we rule out for the flares of AU Mic is driven by high-energy electron beams. Alternatively, high-energy proton beams (consisting of E 10 MeV protons and thus not leaving signatures in the red wing of Ly) could lead to deeper evaporation of cooler material in flares (Sadykov et al., 2024). We plan to investigate this with the RADYN code in future work (McComas et al., 2008).
The prolonged enhancement during the decay phase of several flares, as well as complex flare events such as Flare B, suggests that a more detailed flare model is required to explain an observed persistent heating after high-energy events, which a single heating phase cannot explain during the flare. These advancements need to occur both from modeling the morphology of the flare light curve and in the 1-D radiative-hydrodynamic models. For example, it could be the case that the broadened C II spectral lines during Flare B (Figure 6) may be due to the complex nature of this multi-flare event.
The results of this work are important for understanding flare physics on M dwarfs and for understanding the impact of stellar activity on short-period planets. We require a comprehensive understanding of stellar flare atmospheric physics to be able to model and address whether rocky planets orbiting M dwarfs could retain their atmospheres and be habitable, or not. A thorough understanding of the physics of accelerated non-thermal protons and electrons in energy transport is crucial to this endeavor.
Acknowledgments. ADF acknowledges funding from NASA through the NASA Hubble Fellowship grant HST-HF2-51530.001-A awarded by STScI.
We present the measured energies and equivalent durations in Table 2. We present a recalculation of the white-light (WL) flares for Flares A - M. Additionally, we present the measured flare properties of Flares A - M as observed in the Si III and C III emission lines. We repeat this for the new flares identified in this work (N - T).
Flare | ED | ED | ED | ||||
[s] | [ erg] | [s] | [ erg] | [s] | [ erg] | [s] | |
A | 6388 | — | — | ||||
B | 12531 | ||||||
C | 16935 | ||||||
D | 17985 | ||||||
E | 23049 | — | — | ||||
F | 24819 | ||||||
G | 1740 | ||||||
H | 17515 | — | — | — | — | ||
I | 22993 | — | — | ||||
J | 23473 | ||||||
K | 23653 | ||||||
L | 23983 | ||||||
M | 24493 | 1 | |||||
N | 6090 | ||||||
O | 11622 | ||||||
P | 12616 | ||||||
Q | 16941 | ||||||
R | 17694 | ||||||
S | 19081 | — | — | — | |||
T | 24795 | — | — | — |

### v2

@@
-Probing Non-Thermal Processes in Stellar Flares on AU Mic
+Far-Ultraviolet Emission Line Investigation of Flares on AU Mic
 Abstract
 The role of non-thermal proton energy transportation during solar and stellar flares is largely unknown; a better understanding of this physical process will allow us to rectify longstanding deficiencies in flare models.
-One way to detect the presence of non-thermal protons during flares is through the Orrall-Zirker (OZ) effect whereby an enhanced red wing appears in hydrogen emission lines (e.g., Lyman- at 1215.67 Å).
+One way to detect the presence of non-thermal protons during flares is through the Orrall-Zirker (OZ) effect, proposed by Orrall & Zirker (1976), whereby an enhanced red wing appears in hydrogen emission lines (e.g., Lyman- at 1215.67 Å).
 We analyze archival Hubble Space Telescope/Cosmic Origins Spectrograph G130M (1060 - 1360 Å) observations of the young M dwarf, AU Mic to search for evidence of OZ effect during the impulsive phase of six stellar flares with erg.
@@
 Non-thermal protons are thought to deposit their energy deeper in the atmosphere than non-thermal electrons, thus possibly providing an explanation for the enhancements in the optical continuum, commonly referred to as white-light, that have a photospheric origin (Sadykov et al., 2024; Kerr & Fletcher, 2014; Watanabe et al., 2013; Neidig, 1989; Procházka et al., 2018).
-Non-thermal electrons cannot account for a photospheric origin of white-light flares because they cannot reach deep enough into the atmosphere to cause this heating (Watanabe et al., 2013).
-To determine the cause of white-light flares and other responses to deep energy deposition, we seek an accurate picture of the non-thermal proton acceleration process.
-Constraining the presence of low-energy protons ( MeV) among the accelerated particle population in turn constrains the total energies available to explain the white-light radiation and other flare emission.
+The origin of white light flares is still unknown (Watanabe et al., 2017).
+Non-thermal electrons alone cannot account for a photospheric origin of white-light flares because they cannot penetrate deep enough into the atmosphere to cause the required heating (Watanabe et al., 2013).
+Chromospheric evaporation occurs when non-thermal electrons stream along magnetic field lines and impact the chromosphere.
+Once these non-thermal electrons arrive at the chromosphere, they deposit their energy into the plasma, heating the plasma and increasing pressure.
+The heated plasma is accelerated upwards, resulting in chromospheric evaporation (Fisher et al., 1985a, b, c).
+Observational evidence of chromospheric evaporation includes blue-shifted coronal spectral lines during the impulsive phase of the flare (Milligan & Dennis, 2009; Doschek et al., 1980; Antonucci et al., 1984; Shibata et al., 1995; Graham & Cauzzi, 2015; Polito et al., 2016) and the empirical relation between impulsive and gradual phase emissions known as the Neupert effect (Neupert, 1968; Dennis & Zarro, 1993; Veronig, 2003).
+There are generally thought to be two regimes of chromospheric evaporation: gentle and explosive, whereby the dividing line is a complicated function of the injected electron beam properties; (flux, low-energy cutoff, and power-law index Fisher, 1989; Reep et al., 2015).
+Solar models of chromospheric evaporation tend to favor heating driven by electron beams rather than thermal conduction (Polito et al., 2016).
+In stellar flares, blueshifts have also been observed in hot coronal flare lines (Argiroffi et al., 2019; Inoue et al., 2024), but occasionally they also appear in cooler chromospheric lines such as H (Fuhrmeister et al., 2008; Honda et al., 2018; Notsu et al., 2024; Lu et al., 2025; Kajikiya et al., 2025).
+Blueshifts are typically modeled as filament eruptions or coronal mass ejections in stellar flares.
+Radiative-hydrodynamic models of chromospheric evaporation have been able to produce blueshifts in cool lines (Allred et al., 2005), but the evaporations are too short lived and too slow to explain most of the stellar data.
+More recent radiative-hydrodynamic models suggest that changes in the electron beam can drive deeper material from the chromosphere into the transition region and corona (Kerr et al., 2024; Polito et al., 2023).
+To determine the cause of white-light flares and other responses to deep energy deposition, we seek also an accurate picture of the non-thermal proton acceleration process.
+Constraining the presence of low-energy protons among the accelerated particle population in turn constrains the total energies available to explain the white-light radiation and other flare emission.
 An empirical re-examination of the evidence for these low-energy protons is timely because there has been significant progress in producing power-law spectra of protons extending down to low energies during reconnection (Yin et al., 2024) and in modeling the warm-target transport of low-energy protons from the corona to the chromosphere (Allred et al., 2020).
@@
 Tentative evidence of the OZ effect has been seen in a single star — AU Microscopii (AU Mic; Woodgate et al., 1992).
-AU Mic is a Myr (Bell et al., 2015) M-dwarf (M1Ve; ; Keenan & McNeil, 1989; Plavchan et al., 2020; Gallenne et al., 2022).
+AU Mic is a Myr (Bell et al., 2015) M-dwarf (M1Ve; RR; MM Keenan & McNeil, 1989; Plavchan et al., 2020; Gallenne et al., 2022).
 AU Mic is known for its enhanced stellar activity due to its young age, including a high flaring rate (Feinstein et al., 2022; Gilbert et al., 2022; Tristan et al., 2023), short rotation period ( days; Plavchan et al., 2020), and strong magnetic field (Donati et al., 2023).
@@
 (2023) defined the expected wavelength range to detect the OZ effect as 5.6 - 17.7 Å redward (1380-4365 km s-1) of Ly (see also Canfield & Chang, 1985).
-In 1992, Woodgate et al.
+The OZ effect would be detectable for low energy non-thermal protons with energies ranging from 0.01 to 1 MeV as discussed in Kerr et al.
+(2023) and Canfield & Chang (1985).
+Woodgate et al.
 (1992) presented a tentative detection of the OZ effect on AU Mic through an observed red enhancement in Ly using the Goddard High Resolution Spectrograph on the Hubble Space Telescope (HST).
@@
 To date, this is still the only tentative detection of the OZ effect on another star (Kerr et al., 2023; Robinson et al., 2001).
-This tentative detection highlights that HST could detect the OZ effect here if a sufficient amount of non-thermal protons are accelerated during a flare.
+This tentative detection highlights that HST could detect the OZ effect here if a sufficient amount of non-thermal protons are accelerated during a flare ( erg s-1 cm-2, as modeled in Kerr et al.
+(2023)).
 Feinstein et al.
@@
 In addition to its activity, AU Mic has been a target of observational campaigns due to the presence of two transiting exoplanets (Plavchan et al., 2020; Martioli et al., 2021) and its debris disk (Liu, 2004; Lawson et al., 2023).
-AU Mic b is a Neptune-like exoplanet () on a close orbit, with a semi-major axis au.
-AU Mic c is slightly smaller ( ) and farther than AU Mic b, with a semi-major axis of au (Martioli et al., 2021).
+AU Mic b is a Neptune-like exoplanet (R R⊕) on a close orbit, with a semi-major axis a au.
+AU Mic c is slightly smaller (R R⊕) and farther than AU Mic b, with a semi-major axis of au (Martioli et al., 2021).
 Due to the close proximity of both planets, their evolution has the potential to be shaped by AU Mic’s frequent stellar flares (e.g., Feinstein et al., 2022; do Amaral et al., 2025).
-For example, M-dwarf stars’ frequent and intense flaring may be a source of atmospheric stripping of exoplanets Lammer et al.
-(2007), with exoplanet atmospheric loss being predicted to be partially driven by M-dwarf flaring activity (Neves-Ribeiro et al., 2022).
+For example, M-dwarf stars’ frequent and intense flaring may be a source of atmospheric stripping of exoplanets (Lammer et al., 2007), with exoplanet atmospheric loss being predicted to be partially driven by M-dwarf flaring activity (Neves-Ribeiro et al., 2022).
 The potential of later atmospheric recovery is debated (Tilley et al., 2019).
@@
 Beyond the AU Mic system, a complete understanding of the underlying physics driving magnetic reconnection events and the resulting observable stellar flares will provide a comprehensive understanding of the environments in which most planets evolve.
-M-dwarf stars such as AU Mic are known for their frequent and intense flaring throughout and especially early in their lives (Feinstein et al., 2020b; Crowley et al., 2024).
+M-dwarf stars such as AU Mic are known for their frequent and intense flaring throughout and especially early in their lives (Feinstein et al., 2020; Crowley et al., 2024).
 Stellar CMEs, often accompanied by flares, can potentially be associated with spectral line shifts during a stellar flare, as shown in Ikuta & Shibata (2024).
@@
 Bridging the current gaps in how flares operate will allow an understanding of exoplanet atmospheric responses.
+The models used in Chen et al.
+(2021), where the effects of flares on rocky planets are explored, only use photometric information, so there is no spectroscopic component to these models.
+By neglecting a spectroscopic component, the full impact of the effects of flares on rocky exoplanets are overlooked (Kowalski et al., 2025).
+We do not have many exoplanet spectra that show the direct response to flare irradiation to bridge this gap; thus the models of how flares affect exoplanet atmospheres do not leverage the precision that these spectroscopic observations offer.
 In this work, we present a reanalysis of three HST/COS visits of AU Mic (Feinstein et al., 2022, 2024) to search for evidence of non-thermal proton beams.
@@
 We describe the HST observations, data reduction, and emission line identification in Section 2.
-We identify and characterize new stellar flares in Section 3 and present the first far-ultraviolet (FUV) flare-frequency distribution (FFD) for AU Mic.
+We identify and characterize new stellar flares in Section 3.
 In Section 4, we present the results of our search for evidence of the OZ effect and quantify an identified blue enhancement in several C II and C III emission lines.
@@
 2 Observations
-We observed three visits of AU Mic with HST/COS G130M (Å; Å) under GO 16164 (PI Cauley).
+We observed three visits of AU Mic with HST/COS G130M (Å; Å, Resolving Power, R= ) under GO 16164 (PI Cauley).
 Visit 1 was executed on 2021-05-28 from 03:33:31 — 15:50:43, Visit 2 was executed on 2021-09-23 from 15:06:14 — 22:02:08, and Visit 3 was executed on 2022-10-09 from 12:38:55 — 19:33:40.
@@
 3 The FUV Stellar Flares of AU Mic
+Understanding the time evolution, energy, and rate of flaring for AU Mic is important in understanding specific physical phenomena, such as the OZ effect, observed during flares.
 In this work, we present new stellar flares which occurred during Visit 3.
@@
 Feinstein et al.
-(2022) identified flares in Visits 1 and 2 by identifying outliers in the Si III ( = 1294.55 Å) and C III ( = 1175.95 Å) emission line light curves, presented in Figure 1.
+(2022) identified flares in Visits 1 and 2 by identifying outliers in the Si III ( = 1294.55 Å) and C III ( = 1175.64 Å) emission line light curves, presented in Figure 1.
 Flares were identified by their well-known time-dependent behavior of a sharp initial increase followed by an exponential decay phase.
 We identify an additional seven flares in Visit 3; we follow the same naming convention and introduce flares N - T.
-We uniformly calculated the energies and equivalent durations for flares A - T, where flares A - M occurred in visits 1 and 2. We derive the flare energy using the following prescription:
-where is the distance to the star, is the start time of the flare defined as the time point the flux increases by from the quiescent flux, is the stop time of the flare defined as the time point the flux gets within of the average quiescent flux, , and is the total flux during the defined start and stop times.
-We adopted (Gaia Collaboration et al., 2021).
+We uniformly calculated the energies and equivalent durations for flares A - T, where flares A - M occurred in visits 1 and 2. We derive the flare energy (E) using the following prescription:
+where is the distance to the star, is the start time of the flare defined as the time point the flux increases by 1- from the quiescent flux, is the stop time of the flare defined as the time point the flux gets within 1- of the average quiescent flux, , and is the total flux during the defined start and stop times.
+The 1- fluxes above , are defined as + erg s-1 cm-2 for the white-light flux, + erg s-1 cm-2 for the Si III flux, and + erg s-1 cm-2 for the C III flux.
+We adopted 0.002 pc (Gaia Collaboration et al., 2021).
 To calculate the equivalent duration (ED), first described by Gershberg (1972), we used the following prescription:
-We calculated the energies and durations of the new flares in the “white light” (broadband) wavelength-integrated flux ( Å), Si III wavelength-integrated line flux, and C III wavelength-integrated line flux.
+We calculated the energies and durations of the new flares in the “white light” (broadband) wavelength-integrated flux ( Å), Si III wavelength-integrated line flux ( Å), and C III wavelength-integrated line flux ( Å).
 For the white light flux, we masked the Å due to the lower signal-to-noise at these wavelengths.
@@
 The measured properties of all flares are presented in Table 1.
-3.1 Optical Flares on AU Mic
-AU Mic was observed with the Transiting Exoplanet Survey Satellite, (TESS; Ricker et al., 2015), in Sectors 1 (August 2018) and 27 (July 2020).
-Gilbert et al.
-(2022) and Tristan et al.
-(2025) have both presented independent analyses of the TESS flares on AU Mic.
-We present a reanalysis using a third independent technique.
-We downloaded the TESS 2-minute cadence observations through lightkurve (Lightkurve Collaboration et al., 2018).
-We use the PDC_SAP flux from the SPOC-processed light curves (Jenkins et al., 2010).
-We identify flares using the convolutional neural network (CNN) model stella (Feinstein et al., 2020a, b).
-stella assigns each data point a probability that it is (1) or is not (0) part of a flare.
-We choose this method as the CNN model is able to identify flares without removing rotational photometric variability.
-We considered all flares with an assigned probability .
-We calculated the flare energies and EDs using Equations 1 and 2.
-We identified 54 flares in Sector 1 and 55 flares in Sector 27, for a total of 109 flares in 48.7 days of observations.
-This results in an optical flare rate of 2.24 flares day-1.
-Our sample of TESS flares has erg.
-Our recovered number of flares is slightly higher than but still consistent with previous works (Gilbert et al., 2022; Tristan et al., 2025).
-We note that additional TESS observations of AU Mic are expected in Sector 95 (starting August 2025).
-3.2 FUV and Optical Flare Frequency Distributions
-We compare the FUV and optical flare frequency distributions (FFD) of AU Mic.
-We derive our FUV FFD using the white light HST, Si III, and C III flare energies derived in this work.
-We look at the FFD of multiple spectral lines as this may provide insights into how the emission processes may change over the range of observed flares.
-We derive our optical FFD using observations from the TESS observations.
-All resulting FFDs are presented in Figure 2.
-We fit two distributions to describe the FFD.
-First, we fit a truncated power law of the form
-where is the energy, is a normalization factor, is the slope, is the energy at which the function turns over, is an offset factor, and is the flare rate.
-We fit this form to characterize .
-We use the non-linear least squares model optimization implemented with scipy.optimize.curve_fit to fit for each variable.
-We repeat this process for all FFDs.
-We fit all available energy bins, with the exception of the highest two energy bins in the HST WL and C III FFDs.
-We choose to ignore these bins as the high-energy regime of the FFD is sparsely sampled due to the limited baseline of the HST observations.
-Additionally, we fit a piecewise function of two power-laws to measure the slope, , of the distribution.
-We choose to use this as the descriptor of the slope, rather than in the truncated power-law, as measuring is a more common practice when describing FFDs (e.g., Ilin et al., 2019, 2021).
-We defined our piecewise function as:
-where is the slope and is the y-intercept.
-We present the best-fit values for , , and in Table 1.
-Due to the large uncertainties in for the HST WL and Si III FFDs, we fit these distributions as a single power-law.
-FFD | [erg] | ||
-TESS | |||
-HST | — | ||
-WL | — | ||
-HST | — | ||
-Si III | |||
-HST | |||
-C III |
-We compare the measured TESS FFD to the average FFD of stars of similar ages and to AU Mic presented in Feinstein et al.
-(2024).
-We find that is consistent to within of the FFD slope for stars with K and ages Myr (Feinstein et al., 2024).
-We find that is inconsistent with the measured FFD slope of stars of similar temperatures and age to AU Mic by .
-This discrepancy could be due to the high SNR observations of AU Mic, which result in a more complete FFD, or could be an indication that AU Mic is particularly active for its age.
-A third alternative scenario is that the flares of AU Mic are driven by star-planet interactions with its two short-period gas-giant planets (Ilin & Poppenhaeger, 2022; Feinstein et al., 2024).
-Loyd et al.
-(2018) presented an analysis of inactive and active M dwarfs that were observed as part of the Measurements of the Ultraviolet Spectral Characteristics of Low-mass Exoplanetary Systems (MUSCLES) Treasury survey (France et al., 2016).
-This study included six “inactive” M dwarfs observed with COS/G130M and four “active” M dwarfs STIS/E140M, including AU Mic.
-Loyd et al.
-(2018) measured the FUV flux within specific integrated regions and fit a single power-law with a slope of for active M-dwarfs.
-We note that the inactive M dwarfs had a flare sample size similar to that presented here.
-The shallower FFD slope measured here could indicate that AU Mic is more active than the cumulative sample of M-dwarfs.
-More observations of AU Mic in the FUV are required to properly sample the high-energy regime of the FFD.
 4 Searching Flare-Affiliated for Non-Thermal Processes
@@
 Thus, we calculated spectroscopic light curves of the red wing (1220 - 1233.37 Å) and blue wing (1197.97 - 1206 Å) of the Ly emission line to detect the OZ effect.
-Within the blue wing, we masked S III ( = 1200.97, 1206.53, 1207.51 Å), as identified using the HST/STIS spectrum of AU Mic presented in Rockcliffe et al.
+Within the blue wing, we masked S III ( = 1200.97 Å) and Si III ( = 1206.53, 1207.51 Å), as identified using the HST/STIS spectrum of AU Mic presented in Rockcliffe et al.
 (2023a).
@@
 This analysis was repeated for six of the highest-energy flares in our data set: Flares B, J, K, M, R, and T.
+From the flare properties calculated in Section 3, we can determine that these 6 high-energy flares have Si III energy the Si III energy from the flare identified in Woodgate et al.
+(1992), where the OZ effect was first tentatively detected.
 We present example time-dependent spectra of the Ly wings during Flare B in Figure 2.
@@
 Flare B is a large triple-peaked flare event and the longest-lived event observed in the FUV.
-We present the blue wing, core, and red wing light curves of C II and C III in Figures 4 and 5, respectively.
+We present the blue wing, core, and red wing light curves of C II and C III in Figures 3 and 4, respectively.
 Both figures show that the blue wings of C II and C III appear to respond mainly to the first peak in Flare B, which is further evidence that the second peak in Flare B is a second flare due to its different spectral responses.
@@
 This highlights the unique emission line response for sequential flares.
-This response was analyzed in other C emission lines, but these emission lines lacked the signal-to-noise needed to make a clear observation of the spectral response of the blue and red wings throughout Flare B.
+We searched for evidence of blue wing enhancements across all other emission lines in our observations.
+For emission lines with sufficient signal-to-noise (e.g., N V, Ly wings), we were unable to detect any asymmetrical responses between the blue and red wings of those lines.
+All other emission lines lacked the signal-to-noise to sufficiently model the line profile.
 4.3 Modeling Chromospheric Evaporation
-Chromospheric evaporation occurs when non-thermal electrons stream along magnetic field lines and impact the chromosphere.
-Once these non-thermal electrons arrive at the chromosphere, they deposit their energy into the plasma, heating the plasma and increasing pressure.
-The heated plasma is accelerated upwards, resulting in chromospheric evaporation Fisher et al.
-(1985).
-This phenomenon is observed in the Sun.
-Observational evidence of chromospheric evaporation includes blue-shifted spectral lines during the impulsive phase of the flare (Milligan & Dennis, 2009; Antonucci et al., 1984; Shibata et al., 1995).
 If the blue-enhancement seen in C II and C III is due to chromospheric evaporation, this would imply that the gas emitted at K is being accelerated upwards through the flare loop.
@@
 This model is a higher energy extension of the F11 solar flare model in Allred et al.
-(2005), which produces prominent blueshifts in cool temperature ( K) flare lines.
+(2005), which produces prominent blueshifts in cool temperature (T K) flare lines.
 We modeled individual time steps between the impulsive and decay phases of Flare B using the RH radiative transfer code from Uitenbroek (2001).
-We evaluated the model for Flare B between seconds at s time-intervals.
+We evaluated the model for Flare B between t seconds at t second time-intervals.
 Note that the model timescale is different than the timescale of Flare B, which lasts seconds (1).
@@
 However, none of the rays or the flux account for the discrepancy between the blue and red enhancements of the C II spectral line; there is no significant redshift or blueshift in the model for C II.
+This chromospheric evaporation model does not explain the observed blue enhancement of the C II spectral line, which could be indicative of a lack of chromospheric evaporation during this flare or a deficiency in current models.
 The source of symmetric broadening in non-hydrogenic lines in solar flares is yet not well understood (Kerr, 2023).
@@
 We speculate that the blue enhancement observed here for C III and C II could be explained by emission coming from the expansion of the legs of the filament.
+Solar IRIS observations reveal that C II and Mg II exhibit blue shifts attributed to chromospheric evaporation Li et al.
+(2019).
+Additional analyses of IRIS observations of these spectral lines may provide further insights into spectral tracers of filament eruptions (e.g., Panos et al.
+(2018)).
 This expansion would produce an initial blueshift, followed by a redshift as the filament expands.
@@
 During Flare B, there was an initial similar blue acceleration of C II and C III, in the range of a few hundred km/s.
+Similar models to those used here have been employed for solar flares, and similar line broadening modeling deficiencies have been identified in other spectral lines, such as the Mg II resonance lines in the near UV (Zhu et al., 2019; Hawley et al., 2007).
 Current abilities to model broadening of Mg II for solar flares fall drastically short (Zhu et al., 2019).
@@
 By multiplying by a factor of 30 for the C II resonance lines, the model is better able to describe the broadening throughout the flare; this suggests that an unknown source of atomic line broadening may also explain the C II widths in stellar flares.
-The enhanced symmetric broadening extends far into the wings (Figure 6, s).
+The enhanced symmetric broadening extends far into the wings (Figure 5, t s).
 Determining the physical origin of the missing amounts of the broadening would thus improve the interpretation of asymmetric wings due to mass motions.
@@
 2015 model atom) in the calculations.
-5.1 FUV Flare Fitting
 The flare energies and equivalent durations presented here were derived from the light curve.
@@
 In particular, the model was unable to account for the extended tail of Flare R, which could be due to subsequent sympathetic flares or persistent emission lines during the decay phase.
-We emphasize exercising caution in future work when using this model to calculate flare energy and duration, as it has not consistently provided an accurate fit in this set of flares.
+We emphasize exercising caution in future work when using this model to calculate flare energy (E) and duration, as it has not consistently provided an accurate fit in this set of flares.
 A more flexible model is needed for modeling FUV flares.
@@
 We conclude the following.
-We present the first comparison of the FUV and optical flare frequency distributions (FFDs) of AU Mic (Section 3.2).
-The differences in the FFDs suggest that the average energy budget of the flares dominates in the FUV over the optical, which is consistent with our understanding of flare temperatures.
-We find no evidence of the OZ effect in the high-energy flares analyzed (Section 4); Woodgate et al.
+We find no evidence of the OZ effect in the high-energy flares analyzed (Section 4.1); Woodgate et al.
 (1992) remains the only tentative detection of the OZ effect on another star to-date.
@@
 The results of this study will improve our modeling assumptions of M dwarf flares and may have implications for modeling exoplanet atmospheric responses to stellar flares.
-6.1 Future Work
 Additional archival and future observations of AU Mic are important for understanding the occurrence of the OZ effect (e.g., HST GO 17613).
@@
 Acknowledgments. ADF acknowledges funding from NASA through the NASA Hubble Fellowship grant HST-HF2-51530.001-A awarded by STScI.
+The HST data presented in this article was obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute.
+The specific observations analyzed can be accessed via https://doi.org/10.17909/bh63-4j34 (catalog doi: 10.17909/bh63-4j34).
 We present the measured energies and equivalent durations in Table 1.
@@
 We repeat this for the new flares identified in this work (N - T).
-Flare | ED | ED | ED | ||||
+| Flare | tpeak | E | ED | E | ED | E | ED |
 | [s] | [ erg] | [s] | [ erg] | [s] | [ erg] | [s] | |